OpenStudy (anonymous):

ax + 2xy = 14 for y SHOW WORK PLEASE

OpenStudy (anonymous):

ax+2xy=14 x(a+2y)=14 a-2y=14/x -2y=14/x-a y=\[y=\frac{14}{x}-a \div-2\] Unsure..